OpenStudy (anonymous):

Travis is riding the Ferris wheel at the amusement park. His height can be modeled by the equation H(t) = 22 cospi over 13t + 28, where H represents the height of the person above the ground in feet at t seconds. Part 1: How far above the ground is Travis before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Travis begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Travis' height above the ground reaches a minimum? @jhonyy9 @petiteme @dan815 @asnaseer

OpenStudy (anonymous):

i need help on 2 and 3

OpenStudy (petiteme):

For Part 2: period=2π/B=2π/(π/13)=26 sec

OpenStudy (anonymous):

where does the B come from?

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

wait why do you divide pi/13 by 2pi?

OpenStudy (anonymous):

for 2

OpenStudy (anonymous):

@petiteme for two why did you do that?

OpenStudy (anonymous):

You divide (2pi)/(pi/13) not the other way around. The B value is 2pi/period. You are confining one cycle of the cosine wave to a set about of time. Try playing with some graphing software, like geogebra and others. f(x)=Acos(Bx+C) +D is the equation to play with.

OpenStudy (anonymous):

The B value is just a ratio. It is the ratio between one cycle of a cosine wave and the amount of time to produce one cycle. one cycle/the amount of time to produce one cycle = B value. One cycle for a cosine wave is 2pi. The amount of time to produce one cycle in your equation is 26 seconds. So 2pi/26 = pi/13 which is the B value in your equation.